100.0 mL of 0.7500 M sulfuric acid is titrated with 50.00 mL NaOH. What is the concentration of NaOH?
Answer #1

2NaOH + H2SO4 Na2SO4 + H2O

Moles of H2SO4 = molarity×volume/1000

= 0.7500×100/1000

= 0.075 moles

According to the reaction,

Moles of NaOH = 2×moles of H2SO4

= 2×0.075

= 0.15 moles

Concentration of NaOH = moles×1000/volume

=0.15×1000/50

= 3M

please give feedback!